# references.bib

### v4

@@ 2.5 Baselines
The first baseline is random search of dynamic symbolic policies, which randomly samples a set of trees for each policy. The policies have the same model structure as the dynamic policies (Fig. 2c). The random search policies have the same number of trees in the latent state and are built from the same set of functions and leaf nodes as used in GP. The maximum depth of trees found with random search is fixed to seven. Random search is included as a baseline to demonstrate if GP can find better solutions with less computation than random sampling, therefore showing the additional value of the genetic operators during optimisation.
@@
with $\mathbf{v}=(\mathbf{a},\mathbf{x}^{\ast},1)$ and $\mathbf{B}\in\mathbb{R}^{C\times|\mathbf{v}|}$. In this case, model parameters are optimised using covariance matrix adaption evolution strategies [hansen2016cma]. In all experiments, the dimension of $\mathbf{a}$ was set to five for the NDE baseline. If policies discovered with GP can obtain similar fitness scores as the black-box NDE, the symbolic policies are said to be high-performing.
@@ 2.6 Environments
Our method and the baselines are evaluated on three different environments, as shown in Fig. 3. The first environment is the stochastic harmonic oscillator (SHO), which is a linear environment that can be solved with a linear controller. The policies are tested in a noisy setting, a partial observable setting where only a subset of the state variables is observed, and a setting where the environmental parameters vary. Here, an analytically optimal control policy, linear quadratic Gaussian (LQG) control, is used as an optimal baseline for comparison. The second environment is the acrobot swing up task with bounded continuous action space. This is a non-linear environment that requires a non-linear controller to perform the swing up. The policies solve the acrobot under noise and partial state observability. The third environment is the continuous stirred tank reactor (CSTR). The CSTR is a non-linear industrial system, which shows how well GP can find symbolic policies in an industrial setting. The parameters of the CSTR are varied in every trial and the state is partially observed. The chosen hyperparameter setting of the GP algorithm for each experiment are presented in Table 2 (Appendix 2.4). The choice of the function sets are described in the following subsections.
@@ 2.6.1 Stochastic harmonic oscillator
The SHO is a linear model that describes the position and velocity of an oscillating mass in a single dimension under random perturbations. The SHO is visualised in Fig. 3a and the dynamics of the SHO are given by
$d\mathbf{x}=(\mathbf{Ax}+\mathbf{b}{u})dt+\mathbf{v}d{w}$ (8)
with
with $\mathbf{Q}=\textrm{diag}(0.5,0)$ and $r=0.5$. Here, $\mathbf{Q}$ punishes the policy for the distance between $\mathbf{x}$ and some target state $\mathbf{x}^{\ast}$ and $r$ penalises the control strength. The initial condition $\mathbf{x}_{0}$ is randomly sampled from a normal distribution with zero mean and variance diag(3, 1) for the position and velocity respectively. The target position is chosen randomly by sampling uniformly between -3 and 3 in each trial, while the target velocity is always set to zero. As the optimal LQG control is linear, the function set includes basic arithmetic operators.
@@ 2.6.2 Acrobot swing up task
The acrobot is defined by two connected links, which is visualised in Fig. 3b. The first link is connected to a fixed position. In the swing up task, the goal of the policy is to bring the end point of the second link above a certain threshold height. The dynamics of the acrobot are defined as follows [sutton1995generalization]. Let $\theta_{1}$ and $\theta_{2}$ denote the angles of the first and second link, respectively. Their dynamics are given by
$\ddot{\theta}_{1}$ $=-\frac{d_{2}\ddot{\theta}_{2}+\phi_{1}}{d_{1}}$
$\ddot{\theta}_{2}$ $=\frac{u_{1}+d_{1}^{-1}d_{2}\phi_{1}-m_{2}l_{1}lc_{2}\dot{\theta}_{1}^{2}\sin(\theta_{2})-\phi_{2}}{m_{2}l_{c2}^{2}+I_{2}-d_{1}^{-1}d_{2}^{2}}$
with parameters as defined in Appendix B. In the acrobot experiment that requires two control forces, the state equation of $\dot{\theta}_{1}$ is adapted to $\ddot{\theta}_{1}=({u_{2}-d_{2}\ddot{\theta}_{2}-\phi_{1}})/{d_{1}}$. In practice, we convert this system to a stochastic first-order system
$d\mathbf{x}=f(\mathbf{x},\mathbf{u})dt+\mathbf{V}d\mathbf{w}$
In every experiment, $\mathbf{R}\in\mathbb{R}^{C\times C}$ is set to $0.01\mathbf{I}$, $x^{\ast}$ to 1.5 and the initial states are sampled from $U(-0.1,0.1)$. Observations of $\theta_{i}$ are wrapped between [-$\pi$, $\pi$] and a trial terminates when $\dot{\theta_{1}}>8\pi$ or $\dot{\theta_{2}}>18\pi$. Policies are discovered to solve the acrobot in three different experiments. In the first experiment the state of the acrobot is observable with noise and control force is only applied to the second link. In the second experiment the acrobot state is made partially observable, as the policies no longer observe the angular velocities. In the third experiment, the policies can apply independent forces to both links, and the angular velocities are observed again. The observations of the acrobot include angles, therefore the sine and cosine functions are added to the function set of the GP algorithm, besides the arithmetic operators. As for the SHO, the parameters of the readout function are set to $\mathbf{D}=\mathbf{I}$ and $\mathbf{\Sigma}=0.3\mathbf{I}$ in case of observation noise. In experiment with partial state observability, $\mathbf{D}$ is again changed to exclude the unobserved variables and $\mathbf{\Sigma}$ is shrunken to the correct dimension.
@@ 2.6.3 Continuous stirred tank reactor
The CSTR describes a chemical process following non-linear dynamics [uppal1974dynamic]. The CSTR consists of a main tank that contains a mixture of chemicals and a cooling jacket around the main tank. A visualisation of the CSTR is presented in Fig. 3c. To optimise the chemical process inside the main tank, the temperature has to be stabilised around a setpoint temperature. The temperature in the cooling jacket can be influenced directly by altering the input flow of coolant, which in turn affects the temperature in the main tank. The dynamics of the CSTR are defined as in [antonelli2003continuous], extended with a Wiener process to incorporate stochasticity. Let $\mathbf{x}=(c,T_{r},T_{c})^{\top}$. The CSTR is defined as a stochastic first-order system
$d\mathbf{x}$ $=f(\mathbf{x},\mathbf{u})dt+\mathbf{V}d\mathbf{w}$
with $\mathbf{V}=\textrm{diag}(0.025,3,3)$. The state equation is defined as $f(\mathbf{x},\mathbf{u})=[f_{1}(\mathbf{x},\mathbf{u}),f_{2}(\mathbf{x},\mathbf{u}),f_{3}(\mathbf{x},\mathbf{u})]^{\top}$ with
$f_{1}(\mathbf{x},\mathbf{u})$ $=\frac{q_{r}}{V_{r}}\left(c_{f}-c\right)+k(T_{r})c$
$f_{2}(\mathbf{x},\mathbf{u})$ $=\frac{q_{r}}{V_{r}}\left(T_{f}-T_{r}\right)+\frac{-\Delta H}{\rho C_{p}}k(T_{r})c+\frac{UA}{\rho C_{p}V_{r}}\left(T_{c}-T_{r}\right)$
$f_{3}(\mathbf{x},\mathbf{u})$ $=\frac{q_{c}}{V_{c}}\left(T_{cf}-T_{c}\right)+\frac{UA}{\rho C_{p}V_{r}}\left(T_{r}-T_{c}\right)$
and the parameters as defined in Table 4.
where
The policy has to minimise the distance between the temperature in the reactor and the setpoint temperature $T_{r}^{\ast}$, while also minimising the control, i.e. the coolant flowrate. Only a single experiment is performed regarding the CSTR. This experiment tests policies for their generalisability to an industrial environment. The parameters of the CSTR are different in each trial, to simulate as if the policy is operating in multiple scenarios. In total, eight parameters vary between trials (Table 4). Additionally, the concentration in the reactor is not observed, and the observations of the temperature in the reactor and cooling jacket are noisy. The function set of the GP algorithm is extended to include the exponential and logarithmic functions to aid expressions to deal with the values of the temperatures, which typically range between 200 and 600. To generate the observations, $\mathbf{D}$ is defined such that the concentration is excluded from the observations and $\mathbf{\Sigma}=7.5\mathbf{I}$.
@@ 3 Results
Figure 4: Genetic programming evolves high-performing dynamic symbolic policies. (a) Evolution results of the experiment on the stochastic harmonic oscillator (SHO) with observation noise. The left plot shows the best fitness at every generation during evolution, averaged over 20 independent runs. The methods included are genetic programming for static (GP-S) and dynamic (GP-D) policies, random search (RS), neural differential equation (NDE) and the analytical optimal linear quadratic Gaussian (LQG). The right plot presents the validation fitness obtained on unseen conditions with the best policy evolved in every evolutionary run. (b) Evolution results of the experiment on the acrobot with noisy observations. (c) Evolution results for the experiment in which the state of the SHO is partially observable. (d) Evolution results of the experiment on the acrobot under partial state observability. (e) Evolution results of the experiment on the SHO with varying parameters. (f) Evolution results of the experiment on the acrobot with two control inputs. (g) Evolution results of the experiment on the continuous stirred tank reactor (CSTR) with partial state observability and varying parameters. (h) Best fitness of of each method on the validation set in each experiment. The bold fitness indicates a method is the best for an experiment.
@@ 3.3 Symbolic policies generalise to varying environments
Besides partial observations, realistic settings also require policies to generalise to varying and changing environments. To validate that generalising symbolic policies can be evolved, we test policies in the SHO with parameters that vary between trials (Fig. 4e). In this experiment, the upper bound fitness is again obtained with the LQG controller computed independently for each parameter setting (black curves). The NDE converges close to the upper bound fitness and is consistent across different runs (green curve and dots).
Static policies are first evolved to validate their generalisability (blue curve). GP-S does not converge near the upper bound fitness and the NDE. The validation fitness of the best static policies shows that a ceiling has been reached, as every policy is located close to the average (blue dots). GP-D converges to a similar fitness as GP-S (orange curve). However, the validation fitness of the dynamic symbolic policies shows an interesting pattern (orange dots). Many dynamic symbolic policies obtain a validation fitness that is better than any static policy, and some even compare with NDE policies. However, some dynamic symbolic policies obtain a validation fitness worse than any static policy.
The inconsistent evolution of dynamic symbolic policies is explained by the large search space. When dynamic policies are evolved successfully, they perform better than static policies. The memory allows dynamic policies to discover useful properties that increase generalisability. In our results, the majority of evolutionary runs resulted in dynamic policies that exceed the performance of static policies. Therefore the addition of the latent state in symbolic policies proves to be more beneficial than disadvantageous. Nonetheless, our approach does not consistently produce policies that perform comparably as the NDE.
@@ 3.4 Genetic programming evolves higher-dimensional policies
As the search space of symbolic policies is large, especially for the dynamic symbolic policies, it is important to validate whether high-performing are optimised efficiently when the control dimension increases. In this experiment, the policies can apply control forces to both links to accomplish the acrobot swing-up (Fig. 4f). The NDE converges to a slightly higher fitness during evolution than in the other experiments on the acrobot, because more force can be applied to solve the swing up (green curve). The best NDE policies obtain consistent validation fitness (green dots).
@@ 3.5 Generalising symbolic policies are discovered for an industrial application
In the final experiment, we apply our method to the continuous stirred tank reactor (CSTR), an industrial problem. The temperature in the reactor has to be stabilised at a setpoint temperature by controlling the flow of coolant into the cooling jacket (Fig. 4g). To increase the complexity of the problem, the environment setting is varied between trials and the policies receive partial state observations.
The NDE has not converged after the dedicated number of generations, while being stuck at a poor fitness for the first half of the evolution (green curve). A potential reason for the limited convergence of the NDE could be that they cannot easily approximate the non-linearity required to stabilise the reactor efficiently. The validation fitness of the NDE policies is lower on average than the symbolic policies, and there are a few bad-performing policies (green dots).
GP-S converges to a higher fitness than NDE (blue curve), and the validation fitness shows high consistency among the policies (blue dots). GP-D obtains a better fitness than GP-S during evolution (orange curve). GP-D discovered several policies that outperform GP-S and NDE in terms of validation fitness (orange dots). However, the high-performing dynamic symbolic policies are again not consistently found, as some policies obtain a validation fitness significantly worse than any static policy.
@@
Table 1: The best static and dynamic symbolic policies discovered with genetic programming for every experiment. $u$ represents the function that outputs the control signal. In the dynamic policies, the functions of $a_{i}$ represent the state equations of the latent variables. When a latent variable is not referred to in any of the other expressions, it is labelled as inactive.
Environment Experiment Static policy Dynamic policy
SHO Observation noise $\begin{array}[]{lll}u&=&-0.61y_{2}+x^{\ast}\end{array}$ $\begin{array}[]{lll}u&=&-2a_{1}+2.60a_{2}+x^{\ast}\\ \dot{a}_{1}&=&y_{2}\\ \dot{a}_{2}&=&-u+x^{\ast}\end{array}$
SHO Partial state observability $\begin{array}[]{lll}u&=&0.75x^{\ast}-0.11\end{array}$ $\begin{array}[]{lll}u&=&0.45(a_{1}+x^{\ast})\\ \dot{a}_{1}&=&2a_{2}-u+x^{\ast}\\ \dot{a}_{2}&=&-a_{2}-0.99u+y_{1}\end{array}$
SHO Varying parameters $\begin{array}[]{lll}u&=&-1.10y_{1}-0.73y_{2}\\ &&+1.83x^{\ast}+0.17\end{array}$ $\begin{array}[]{lll}u&=&-a_{1}^{3}-a_{2}+x^{\ast}\\ \dot{a}_{1}&=&27.76(-a_{1}+y_{1}+y_{2}-x^{\ast})\\ \dot{a}_{2}&=&0.32a_{1}\end{array}$
Acrobot Observation noise $\begin{array}[]{lll}u&=&-y_{3}+1.29\sin(y_{4})\end{array}$ $\begin{array}[]{lll}u&=&2a_{2}-\cos(2a_{1})\\ \dot{a}_{1}&=&2y_{1}-2y_{2}\\ \dot{a}_{2}&=&5.47\sin(\sin(y_{1}))\end{array}$
Acrobot Partial state observability No successful policies $\begin{array}[]{lll}u&=&1.87a_{1}+\cos(a_{2})\\ \dot{a}_{1}&=&2.06y_{1}\\ \dot{a}_{2}&=&y_{1}-2.68\end{array}$
CSTR Standard $\begin{array}[]{lll}u&=&T_{c}^{4}(T_{r}^{\ast})^{-3}-T_{r}\end{array}$ $\begin{array}[]{lll}u&=&a_{1}^{2}a_{2}+2a_{1}\log(T_{r}^{\ast})\\ \dot{a}_{1}&=&(2.92-a_{1})(u_{1}+2.98)\\ \dot{a}_{2}&=&(u_{1}-a_{2}+2.37)(a_{2}-T_{r}^{\ast}+T_{r})\end{array}$
Figure 5: Dynamic symbolic policies are transparent. (a) Simulated trajectory of the stochastic harmonic oscillator (SHO) controlled by a dynamic symbolic policy. The parameters of the oscillator change over time, as is displayed in the top row. The plot in the second row shows the state and control in the trajectory. The shaded area around the states is observed by the policy. For each environment, the dashed line indicates the setpoint. The latent state variables of the policy are presented in the bottom row. The policy used in this simulation is described in Table 1 (SHO, Varying parameters, Dynamic) (b) A simulated trajectory of the acrobot with observation noise, where a dynamic symbolic policies applies force the second link. The top row shows the position of the two links over time. The second row presents the control force, and the bottom row shows the latent state variables. The policy is described in Table 1 (Acrobot, Observation noise, Dynamic). (c) Simulated trajectory of the CSTR controlled with a dynamic symbolic policy. The top row shows the temperature of the reactor and cooling jacket when the policy controls the coolant inflow with varying setpoint temperatures. The second row presents the concentration of the reactant in the reactor. The bottom row shows the behavior of the latent state of the policy. The policy used in this simulation is described in Table 1 (CSTR, Dynamic).
3.6 The latent states in a dynamic symbolic policy are transparent

### v5

@@ 3.5 Baselines
The first baseline is random search of dynamic symbolic policies, which randomly samples a set of trees for each policy. The policies have the same model structure as the dynamic policies (Fig. 2c), and consist of the same set function and leaf nodes. The maximum depth of trees is fixed to seven. Random search is included as a baseline to demonstrate if GP can find better solutions with less computation than random sampling, therefore showing the additional value of the genetic operations during optimisation.
@@
with $\mathbf{v}=(\mathbf{a},\mathbf{x}^{\ast},1)$ and $\mathbf{B}\in\mathbb{R}^{C\times|\mathbf{v}|}$.
In this case, model parameters are optimised using covariance matrix adaption evolution strategies [hansen2016cma]. In all experiments, the dimension of $\mathbf{a}$ was set to five for the NDE baseline. The NDE policies are included as a baseline to investigate whether symbolic policies can compete with a proven black-box method.
@@ 3.6 Environments
Our method and the baselines are evaluated on three different environments, as shown in Fig. 3. The policies are tested on a linear and non-linear classical control task and on a non-linear industrial problem. The chosen hyperparameter setting of the methods for each experiment are presented in Table 3 (Appendix 3.4). The choice of the function sets are described in the following subsections.
@@ 3.6.1 Stochastic harmonic oscillator
The stochastic harmonic oscillator (SHO) is a linear system that describes the position and velocity of an oscillating mass in a single dimension under random perturbations. The SHO is visualised in Fig. 3a. The goal of the controller is to stabilise the mass at a target position while minimising the control usage. The SHO is easily solved with a linear controller, therefore we compute the linear quadratic Gaussian (LQG) control as an optimal baseline for comparison. The policies are tested in a setting with noisy observations, a setting where only the position of the oscillator is observed, and a setting where the environmental parameters vary per trajectory. To discover generalising policies, each policy is tested on a batch of trajectories, with random initial conditions and targets. Since the optimal controller is linear, the function set of GP consists of the basic arithmetic operators. See Appendix B for details on the dynamics, the fitness function and the specific definitions of the experiments.
@@ 3.6.2 Acrobot swing up task
The acrobot is defined by two connected links, where the first link is connected to a fixed position (Fig. 3b). The goal is to swing the end point of the second link above a threshold, while minimising the applied force. The dynamics of the acrobot are non-linear, and to accomplish the swing up a non-linear controller is required. The policies solve the acrobot in experiments with noisy observations including and excluding the observations velocity, and a third experiment is conducted where control can be applied to both links individually. The observations of the acrobot include angles, therefore the sine and cosine functions are added to the function set of the GP algorithm, besides the arithmetic operators. The dynamics and the fitness function are described in Appendix C.
@@ 3.6.3 Continuous stirred tank reactor
The continuous stirred tank reactor (CSTR) describes a chemical process that is governed by non-linear dynamics [uppal1974dynamic]. The CSTR consists of a main tank that contains a mixture of chemicals and a cooling jacket around the main tank. A visualisation of the CSTR is presented in Fig. 3c. To optimise the chemical process inside the main tank, the temperature has to be stabilised around a setpoint temperature. The temperature in the cooling jacket can be influenced directly by altering the input flow of coolant, which in turn affects the temperature in the main tank. in our experiment, the parameters of the CSTR are varied in every trial (Table III). The policy receives noisy observations of the temperature in the reactor and cooling jacket, but the concentration is not observed. The function set of the GP algorithm is extended to include the exponential and logarithmic functions to aid expressions to deal with the values of the temperatures, which typically range between 200 and 600. The dynamics and the values of the environment’s parameters are detailed in Appendix D, as well as the fitness function.
@@ 4 Results
Figure 4: Genetic programming evolves high-performing dynamic symbolic policies. (a) Evolution results of the experiment on the stochastic harmonic oscillator (SHO) with observation noise. The left plot shows the best fitness at every generation during evolution, averaged over 20 independent runs. The methods included are genetic programming for static (GP-S) and dynamic (GP-D) policies, random search (RS), neural differential equation (NDE) and the analytical optimal linear quadratic Gaussian (LQG). The right plot presents the validation fitness obtained on unseen conditions with the best policy evolved in every evolutionary run. (b) Evolution results of the experiment on the acrobot with noisy observations. The LQG is excluded because it cannot be computed given non-linear dynamics. (c) Evolution results for the experiment in which the state of the SHO is partially observable. (d) Evolution results of the experiment on the acrobot under partial state observability. (e) Evolution results of the experiment on the SHO with varying parameters. (f) Evolution results of the experiment on the acrobot with two control inputs.
@@ 4.2 Dynamic symbolic policies remain robust under partial state observability
Table 1: The fitness of the best policies on validation trajectories for every experiment. The complexity of the best policies is also displayed as the size. The methods include random search (RS), neural differential equation (NDE) and genetic programming for static (GP-S) and dynamic (GP-D) policies. The bold fitness indicates that a method obtained the best fitness in an experiment. The complexity is computed by summing the number of mathematical operations, variables and constants in the policies.
RS NDE GP-S GP-D
Environment Experiment Fitness Size Fitness Size Fitness Size Fitness Size
SHO Observation noise -56.11 15 -48.85 215 -57.28 5 -49.66 19
SHO Partial state observability - -72.76 195 -903.44 5 -63.35 21
SHO Varying parameters - -47.76 215 -110.25 13 -45.56 25
Acrobot Observation noise -123.64 6 -98.45 231 -115.31 8 -112.82 20
Acrobot Partial state observability - -101.47 191 -250.0 1 -110.23 12
Acrobot Two control inputs - -97.02 272 -103.43 12 -100.69 18
CSTR Standard - -1017.72 215 -1059.31 9 -845.33 27
@@ 4.3 Symbolic policies generalise to varying environments
Besides partial observations, realistic settings also require policies to generalise to varying and changing environments. To validate that generalising symbolic policies can be evolved, we test policies in the SHO with parameters that vary between trials (Fig. 4e). In this experiment, the upper bound fitness is again obtained with the LQG controller that is computed independently for each parameter setting (black curves). The NDE converges close to the upper bound fitness and is consistent across different runs (green curve and dots).
Static policies are evolved to validate their generalisability (blue curve). GP-S does not converge near the upper bound fitness and the NDE. The validation fitness of the best static policies shows that a ceiling has been reached, as every policy is located close to the average (blue dots). GP-D converges to a similar fitness as GP-S (orange curve). However, the validation fitness of the dynamic symbolic policies shows an interesting pattern (orange dots). Many dynamic symbolic policies obtain a validation fitness that is better than any static policy, and some even compare with NDE policies. However, some dynamic symbolic policies obtain a validation fitness worse than any static policy.
The inconsistent evolution of dynamic symbolic policies is explained by the large search space. When dynamic policies are evolved successfully, they perform better than static policies. The memory allows dynamic policies to discover useful properties that increase generalisability. In our results, the majority of evolutionary runs resulted in dynamic policies that exceed the performance of static policies. Therefore the addition of the latent state in symbolic policies proves to be more beneficial on average. Our approach does not consistently produce policies that perform comparably as the NDE policies, though this comes at the expense of obtaining white-box solutions.
Table 2: The best static and dynamic symbolic policies discovered with genetic programming for every experiment. $u_{j}$ represents the function that outputs the control signal. In the dynamic policies, the functions of $a_{i}$ represent the state equations of the latent variables. When a latent variable is not referred to in any of the other expressions, it is labelled as inactive.
Environment Experiment Static policy Dynamic policy
SHO Observation noise $\begin{array}[]{lll}u_{1}&=&-0.61y_{2}+x^{\ast}\end{array}$ $\begin{array}[]{lll}u_{1}&=&-2a_{1}+2.60a_{2}+x^{\ast}\\ \dot{a}_{1}&=&y_{2}\\ \dot{a}_{2}&=&-u_{1}+x^{\ast}\end{array}$
SHO Partial state observability $\begin{array}[]{lll}u_{1}&=&0.75x^{\ast}-0.11\end{array}$ $\begin{array}[]{lll}u_{1}&=&0.45(a_{1}+x^{\ast})\\ \dot{a}_{1}&=&2a_{2}-u_{1}+x^{\ast}\\ \dot{a}_{2}&=&-a_{2}-0.99u_{1}+y_{1}\end{array}$
SHO Varying parameters $\begin{array}[]{lll}u_{1}&=&-1.10y_{1}-0.73y_{2}\\ &&+1.83x^{\ast}+0.17\end{array}$ $\begin{array}[]{lll}u_{1}&=&-a_{1}^{3}-a_{2}+x^{\ast}\\ \dot{a}_{1}&=&27.76(-a_{1}+y_{1}+y_{2}-x^{\ast})\\ \dot{a}_{2}&=&0.32a_{1}\end{array}$
Acrobot Observation noise $\begin{array}[]{lll}u_{1}&=&-y_{3}+1.29\sin(y_{4})\end{array}$ $\begin{array}[]{lll}u_{1}&=&2a_{2}-\cos(2a_{1})\\ \dot{a}_{1}&=&2y_{1}-2y_{2}\\ \dot{a}_{2}&=&5.47\sin(\sin(y_{1}))\end{array}$
Acrobot Partial state observability No successful policies $\begin{array}[]{lll}u_{1}&=&1.87a_{1}+\cos(a_{2})\\ \dot{a}_{1}&=&2.06y_{1}\\ \dot{a}_{2}&=&y_{1}-2.68\end{array}$
CSTR Standard $\begin{array}[]{lll}u_{1}&=&T_{c}^{4}(T_{r}^{\ast})^{-3}-T_{r}\end{array}$ $\begin{array}[]{lll}u_{1}&=&a_{1}^{2}a_{2}+2a_{1}\log(T_{r}^{\ast})\\ \dot{a}_{1}&=&(2.92-a_{1})(u_{1}+2.98)\\ \dot{a}_{2}&=&(u_{1}-a_{2}+2.37)(a_{2}-T_{r}^{\ast}+T_{r})\end{array}$
@@ 4.4 Genetic programming evolves higher-dimensional policies
As the search space of symbolic policies is large, especially for the dynamic symbolic policies, it is important to validate whether high-performing policies are found efficiently when the control dimension increases. In this experiment, the policies can apply control forces to both links to accomplish the acrobot swing-up (Fig. 4f). The NDE converges to a slightly higher fitness during evolution than in the other experiments on the acrobot, because more force can be applied to solve the swing up (green curve). The best NDE policies obtain consistent validation fitness (green dots).
@@ 4.5 Generalising symbolic policies are discovered for an industrial application
In the final experiment, we apply our method to the continuous stirred tank reactor (CSTR) as an industrial problem with non-linear dynamics. The temperature in the reactor has to be stabilised at a setpoint temperature by controlling the flow of coolant into the cooling jacket (Fig. 5a). To increase the complexity of the problem, the environment setting is varied between trials and the policies receive noisy partial state observations.
The NDE has not converged after the dedicated number of generations, while being stuck at a poor fitness for the first half of the evolution (green curve). A potential reason for the limited convergence of NDEs could be that they cannot easily approximate the non-linearity required to stabilise the reactor efficiently. The validation fitness of the NDE policies is on average lower than the symbolic policies, and there are a few bad-performing policies (green dots).
GP-S converges to a higher fitness than NDE (blue curve), and the validation fitness shows high consistency among the policies (blue dots). GP-D obtains a better fitness than GP-S during evolution (orange curve). GP-D discovered several policies that outperform GP-S and NDE in terms of validation fitness (orange dots). However, the high-performing dynamic symbolic policies are not consistently found, as some policies obtain a validation fitness significantly worse than any static policy.
@@
Figure 5: Dynamic symbolic policies generalise to industrial control settings. (a) Evolution results of the experiment on the continuous stirred tank reactor (CSTR) with partial state observability and varying parameters. The left plot shows the best fitness at every generation during evolution, averaged over 20 independent runs. The methods included are genetic programming for static (GP-S) and dynamic (GP-D) policies and neural differential equation (NDE). The right plot presents the validation fitness obtained on unseen conditions with the best policy evolved in every evolutionary run. (b) Simulated trajectory of the CSTR controlled with a dynamic symbolic policy. The top row shows the temperature in the reactor and cooling jacket and the controlled coolant flowrate into the cooling jacket with varying setpoint temperatures. The bottom row shows the latent variables of the policy. The policy used in this simulation is described in Table 2 (CSTR, Dynamic). (c) The temperature of the reactant that flows into the main reactor increases throughout the trajectory. (d) The observations of the temperature are made more noisy. (e) The dynamics of the temperature in the main reactor are influenced by an external sinusoidal wave, simulating the change of temperature outside of the reactor during the day and night.
4.6 The latent variables in dynamic symbolic policies are transparent